# SymBoltz.jl: a symbolic-numeric, approximation-free and differentiable linear Einstein-Boltzmann solver

### v1

@@ 3.3 Computing power spectra
using Unitful, UnitfulAstro # for Mpc unitks = 10 .^ range(-4, 1, length=200) / u"Mpc"Ps = spectrum_matter(prob, ks)ls = 10:10:2000Cls = spectrum_cmb([:TT, :TE, :EE], prob, ls)Calling spectrum_matter and spectrum_cmb with prob will automatically select a grid of $k$ to solve the perturbations for, and interpolate between them when computing the spectra. This is a common interpolation trick in Boltzmann solvers to integrate fewer perturbation modes. The functions can be called with the solution sol instead, but will then interpolate only between the $k$ that sol was solved for, leaving it to the user to ensure sufficient sampling density. Figure 3 shows that the spectra agree with CLASS to 11‰-11% or better for the chosen parameters.
@@ 3.4 Differentiable Fisher forecasting
Fisher forecasting is a powerful technique for predicting how strong parameter constraints that can be placed by data with given errors. It requires accurate derivatives and is a nice application for automatic differentiation.
Near a peak $\boldsymbol{\theta}_{0}$, where derivatives vanish, a log-likelihood function of parameters $\boldsymbol{\theta}$ is approximated by the Taylor series
$\log L(\boldsymbol{\theta})\approx\log L(\boldsymbol{\theta}_{0})-\sum_{i,j}F_{ij}(\theta_{i}-\bar{\theta}_{i})(\theta_{j}-\bar{\theta}_{j}),$ (9)
where $F$ is the Fisher (information) matrix with elements
$F_{ij}=-\frac{1}{2}\frac{\partial^{2}\log L}{\partial\theta_{i}\,\partial\theta_{j}}\bigg\rvert_{\,\boldsymbol{\theta}=\boldsymbol{\theta}_{0}}.$ (10)
Intuitively, $F$ measures how sharp the peak is, or how sensitive $\log L$ is to changes in different directions in parameter space. Fisher forecasting is powered by the Cramér-Rao bound, which guarantees that $F^{-1}_{ij}$ is a lower bound $\big\lvert C_{ij}\big\rvert\geq\big\lvert F^{-1}_{ij}\big\rvert$ for the covariance $C_{ij}$ between model parameters $\theta_{i}$ and $\theta_{j}$. The inequality is more saturated the better the likelihood approximation (9) is (i.e. the ?more Gaussian? the probability distribution is, for which the expansion is exact). Thus, computing $F_{ij}$ at a peak and inverting it gives the tightest parameter constraints one can hope for.
Since $F_{ij}$ depends on derivatives of $\log L$, Fisher forecasting traditionally involves error-prone finite differences and step size tuning. This problem is avoided with automatic differentiation.
@@
Assuming the measurements of each $\bar{C}_{l}$ are normally distributed and uncorrelated, the log-likelihood for this experiment is $\chi^{2}$:
$\log L(\boldsymbol{\theta})=-\frac{1}{2}\sum_{l}\left(\frac{C_{l}(\boldsymbol{\theta})-\bar{C}_{l}}{\sigma_{l}}\right)^{2}.$ (12)
In this case, the Fisher matrix (10) becomes
$F_{ij}=\sum_{l}\frac{\partial C_{l}}{\partial\theta_{i}}\frac{1}{\sigma_{l}^{2}}\frac{\partial C_{l}}{\partial\theta_{j}}\bigg\rvert_{\,\boldsymbol{\theta}=\boldsymbol{\theta}_{0}},$ (13)
evaluated in some fiducial cosmology $\boldsymbol{\theta}=\boldsymbol{\theta}_{0}$.
The derivatives $\partial C_{l}/\partial\theta_{i}$ are hard to compute and perfect candidates for automatic differentiation with ForwardDiff.jl: {codebox}
using ForwardDiff: jacobianvary = [ M.g.h, M.c.Ω₀, M.b.Ω₀, M.b.YHe, M.ν.Neff, M.h.m_eV, M.X.w₀, M.X.wₐ, M.I.ln_As1e10, M.I.ns,]genprob = parameter_updater(prob, vary)ls, ls′ = 100:1:1000, 100:25:1000Cl(θ) = spectrum_cmb(:TT, genprob(θ), ls, ls′)θ₀ = map(par -> pars[par], vary)dCl_dθ = jacobian(Cl, θ₀)Here vary orders the subset of parameters to differentiate with respect to, and genprob is a function that generates a new problem with updated parameters θ. Then Cl specifies a function that computes $C_{l}(\boldsymbol{\theta})$ for ls, interpolating from a coarser grid ls′. This machinery just converts from SymBoltz’ parameter-to-value mapping to a function of an array of parameter values, as jacobian requires a pure mathematical function $\boldsymbol{f}:\mathbb{R}^{a}\rightarrow\mathbb{R}^{b}$ to differentiate. Finally $\partial C_{l}/\partial\theta_{i}$ is computed with forward-mode automatic differentiation and stored in dCl_dθ, as shown in fig. 4.
Figure 4: Normalized derivatives $(\partial C_{l}/\partial\theta_{i})/C_{l}$ of a CMB $T$ power spectrum with respect to cosmological parameters $\theta_{i}$ from SymBoltz and automatic differentiation (AD; colored lines) versus CLASS and central finite differences (FD; gray dashes). CLASS uses the precision parameters in appendix B and finite differences with $5\%$ relative step sizes.
@@
It is now trivial to compute the Fisher matrix (13). Inverting it forecasts the constraints in fig. 5. They are in excellent agreement with finite difference results from CLASS, but these required significant tuning of precision parameters and step sizes.
@@ 3.5 Differentiable MCMC sampling with supernova data
Finally, we show that SymBoltz can output differentiable results for gradient-based Markov chain Monte Carlo (MCMC) samplers, like the No-U-Turn Sampler (NUTS). This state-of-the-art method uses likelihood gradients to move more intelligently in parameter space than the ?blind? Metropolis-Hastings algorithm.
We predict apparent magnitudes $m(z)$ of Type Ia supernovae as a function of redshift $z$ and fit them to 1048 observations $(z_{i},m_{i})$ from the Pantheon dataset (Scolnic et al., 2018). Such supernovae are standard candles with constant absolute magnitude $M\approx-19.3$. Their apparent magnitudes are $m(z)=M+\mu(z)$, where $\mu(z)=5\lg(d_{L}(z)/10\text{ pc})$ is the distance modulus from the background-derived luminosity distance [^11] This expression is valid for any $\Omega_{k0}$ with complex $\operatorname{sinc}(x)=\sin(x)/x$, but can be split into branches for positive, negative and zero $\Omega_{k0}$.
@@
We define the likelihood by a multivariate normal distribution of the 1048 supernovae with (constant) covariance matrix $C$ given in detail by Scolnic et al. (2018). To compute it, we interface SymBoltz with the probabilistic programming framework Turing.jl (Fjelde et al., 2025), which allows a high-level description of the probabilistic model equivalent to the log-likelihood
$\log L=-\frac{1}{2}\sum_{i,j}\big(m(z_{i})-m_{i}\big)C^{-1}_{ij}\big(m(z_{j})-m_{j}\big).$ (15)
@@
We use the NUTS sampler in Turing. The code for this example can be found in the paper’s notebook. For the flat $w_{0}\text{CDM}$ model ($\Omega_{k0}=0$, $w_{a}=0$, fixed $\Omega_{r0}$) it gives the constraints in fig. 6. The results are not interesting, but show that differentiable parameter inference works with SymBoltz. Performance improvements are needed to make this viable also for perturbation-derived spectra.
@@
Figure 6: Parameter inference on 1048 Type Ia supernovae from the full Pantheon supernova dataset, using 5000 MCMC samples with the gradient-based NUTS sampler in Turing.jl and differentiable theory predictions from SymBoltz.jl.
4 Discussion of design synergies and tradeoffs
A symbolic-numeric interface, lack of approximations and differentiability are interesting features in their own rights. They also constitute a self-reinforcing synergy explained in fig. 7. Like any Boltzmann solver, SymBoltz makes some tradeoffs in its design.
Symbolic-numericinterfaceFree ofapprox-imationsAutomaticdifferentiationDetect sparsity of JJCalculate $J$ robustlyEasy to specify model
Figure 7: A symbolic-numeric interface, approximation-freeness and differentiability are three important features of SymBoltz that form a synergy. For example: automatic differentiation helps to calculate the ODE Jacobian robustly; which is needed by implicit solvers to integrate stiff ODEs without approximations; which makes it easier to write down models in a simple symbolic form; which can be used to detect the sparsity pattern of the Jacobian; which in turn increases the efficiency of the implicit ODE solver. This synergy creates a self-reinforcing design.
4.1 Symbolic vs. numeric interface
As explained in sections 1.1 and 2.1, modifications to most Boltzmann solvers are made by editing the low-level numerical source code directly, while changes to SymBoltz are made in a high-level symbolic interface that abstracts away internal details.
We think three properties make the linearized Einstein-Boltzmann equations particularly attractive for symbolic abstraction. First, on the ?outside?, they are fundamentally one (large) system of equations with a predictable general structure. Second, on the ?inside?, they are complicated to solve in multiple stages due to their structure, stiffness and need for speed. Third, they are heavily subject to modifications as the true cosmological model is unknown. To simplify the modeling process, these aspects make it valuable to abstract the ?inside? from the ?outside?.
Purely numerical codes generally put the programmer first. They are adapted to the computational structure of the problem. Users have full freedom to tweak problem-specific details, such as implementing complex approximation schemes. Without abstraction layers, the code says exactly what it does. This can be worthwhile for custom-tailoring codes to ?workhorse? cosmological models like $\mathrm{\Lambda CDM}$ for maximum performance and demanding MCMC analyses. However, it can result in an overwhelming monolithic design that is hard to read and modify.
The symbolic interface in SymBoltz prioritizes the modeler. Its design matches the physical structure of the model. Manual chores are automated with compile-time logic on the equations. Users can more easily change equations without understanding internals provided that they can be written in a straightforward form that is compatible with the symbolic interface. This is not a major restriction for the Einstein-Boltzmann equations, as they have a plain and predictable structure when linearized and free of approximations. Nevertheless, to bridge this gap, SymBoltz lets users call arbitrary numerical functions that escape the symbolic-to-numeric compilation, if necessary.
This abstraction layer works only if fits the structure of the underlying equations. A key to envision the design was to look past the computational pipeline structure of Boltzmann solvers (e.g. background $\rightarrow$ perturbations $\rightarrow\ldots$), and rather organize the code primarily by physical components like the metric, gravitational theory and particle species. Components are self-contained and joined into full cosmological models. As more models are added, the complexity of this modular design stays constant, while monolithic ?all-in-one? codes become complicated.
This structure can be unfamiliar to those used to the layout of other codes. The modular structure scales best with a growing number of submodels for individual physical components. But SymBoltz also provides a full unstructured $\mathrm{\Lambda CDM}$ model where everything is merged into one large system, which is easier to work with if one wants to modify arbitrary small parts of the equations. This leaves the choice to the user, who can select between the best of both worlds.
Mixing symbolic and numerical computing can be susceptible to performance problems. Numerics should be fast and symbolics is usually slow, so ?perceived performance? suffers if the two are interleaved. But SymBoltz disentangles the expensive symbolic operations in section 2.1 from numerics, so they are performed only before and after performance-critical tasks.
We hope the design of SymBoltz provides a platform for easy exploration of alternative models, regardless of which sector of the equations one is interested in modifying.
4.2 Approximation-freeness vs. performance
As seen in section 1.2, traditional codes reduce stiffness by approximating the equations for explicit integrators, while SymBoltz solves the full stiff system with implicit methods.
Without approximations, SymBoltz needs only one set of equations to solve. These equations are easy to write down in pure symbolic form, which pairs nicely with SymBoltz’ high-level symbolic interface. With approximations, traditional solvers need more complicated infrastructure to switch between versions of the same equations. This fits well to a low-level and fully numerical code that maximizes implementation freedom, like CLASS.
Approximation schemes do not only alleviate stiffness, but also improve performance by reducing the ODE size. However, Lesgourgues & Tram (2011); Hahn et al. (2024) found that the TCA and UFA provide only marginal speedups compared to use of implicit solvers. On the other hand, the RSA can provide significant speedups with high $l_{\text{max}}$ (Lesgourgues & Tram, 2011; Moser et al., 2022). However, more load is put on the modeler to derive, implement and validate multiple approximations. This process must generally be repeated for modified models that can easily reintroduce stiffness or invalidate the approximations.
Implicit solvers take more expensive steps than explicit ones. They solve a nonlinear system for the next unknowns at every time step, which gets costly as the ODE size grows (e.g. higher $l_{\text{max}}$). This is often done with Newton’s method, which iterates over linear matrix solutions $A\boldsymbol{x}=b$, where $A$ involves the ODE Jacobian $J$. But several factors mitigate this slowdown.
First and foremost, implicit solvers take longer and fewer steps due to better stability properties. This is highly solver-dependent, however: two different implicit solvers can perform very differently on one stiff problem. DifferentialEquations.jl lets SymBoltz easily switch between a large suite of implicit solvers. The background (with stiff thermodynamics) is solved robustly with Rodas4P/Rodas5P, and the perturbations efficiently with TRBDF2/KenCarp4/Rodas5P for low/medium/high precision. Notably, CLASS includes and defaults to a custom implicit solver ndf15 and performs great (Lesgourgues & Tram, 2011). In contrast, SymBoltz can choose between a range of compatible solvers to fulfill different performance and precision requirements.
Second, a key optimization for implicit solvers is to compute $J$ efficiently, reuse it over time steps and LU-factorize $A$ to speed up successive linear system solutions until Newton’s method no longer converges due to outdated Jacobians. The TRBDF2 and KenCarp4 solvers do this. When needed, $J$ is updated with the method in section 2.2, which is cheaper than finite differences. The linear matrix solver used in Newton’s method also matters, as they are among the most thoroughly optimized numerical algorithms. DifferentialEquations.jl interoperates with LinearSolve.jl, which lets SymBoltz easily swap the linear solver in the ODE solver. On our hardware we find that performance improves by $5\times$ with a recursive LU factorization algorithm from RecursiveFactorization.jl or Intel’s (proprietary) Math Kernel Library MTK.jl over Julia’s default OpenBLAS backend.
Third, sparse matrix methods can speed up the linear solver as the system grows in size and the fraction of zeros increases. SymBoltz can find the exact sparsity pattern from the symbolic Jacobian in section 2.1.5, and LinearSolve.jl provides several sparse matrix methods. Support for this is still ongoing work.
Although SymBoltz sacrifices some performance to get rid of approximations, these countermeasures make the impact less severe than one might fear. In return, the approximation-free structure is a major simplification and pairs well with the high-level interface for simple symbolic equations. There is still room to improve performance without resorting to approximations by tuning precision tolerances, supporting sparse matrix methods, exploring implicit-explicit (IMEX) solvers (e.g. Kennedy & Carpenter (2003)) and smarter sampling of $k$ and $l$, for example.
Further performance optimizations and comparisons are left for future work. SymBoltz is fast enough for single runs and interactive usage, but not yet for MCMC analyses with perturbation-derived quantities. Of course, SymBoltz can also implement approximation schemes in the future, but its symbolic nature can make this harder than in other codes. As a counterweight to other codes, SymBoltz’ primary mission will be to solve full stiff equations and keep approximations optional and secondary.
4.3 Forward-mode vs. reverse-mode automatic differentiation
At this time SymBoltz is only well-tested with forward-mode automatic differentiation. However, we saw in section 1.3 that reverse-mode is more attractive for applications with more inputs than outputs. This is the case for popular applications with scalar likelihood or loss functions, such as MCMC parameter inference and training neural network emulators. Reverse-mode could be particularly powerful for parameter inference with next-generation surveys, where experimental nuisance parameters must also be sampled in large $\mathcal{O}(100)$-dimensional parameter spaces, even if eventually marginalized over. However, forward-mode has better characteristics when performing sensitivity analyses. For example, computing $\partial P(k;\boldsymbol{\theta})/\partial\theta_{i}$ or $\partial C_{l}(\boldsymbol{\theta})/\partial\theta_{i}$ (e.g. Fisher forecasting in section 3.4) is faster with forward-mode, since one typically wants $\mathcal{O}(100\text{-}1000)$ $k$ or $l$, but only has $\mathcal{O}(10)$ parameters.
Of course, automatic differentiation is just an additional feature. One can run the code with or without automatic differentiation, or use finite differences instead. It does not pose any tradeoff, but robust support for both forward-mode and reverse-mode remains a future goal that would make SymBoltz more powerful for different applications. This situation will improve as the differentiable Julia ecosystem continues to evolve.
5 Conclusion and future potential

### v2

@@ 3.3 Computing power spectra
Table 1: Time to compute the power spectra in fig. 3.
Code (approximations) $P(k)$ $C_{l}$
SymBoltz (no approximations) 0.3 s 3.1 s
CLASS (only tight-coupling approximation) 5.2 s 8.7 s
CLASS (all approximations) 0.5 s 1.6 s
@@
ks = 10 .^ range(-1, 4, length = 100)Ps = spectrum_matter(prob, ks)ls = vcat([2,3,5,10], 20:20:2500)jl = SphericalBesselCache(ls)Cls = spectrum_cmb([:TT, :EE, :TE], prob, jl)The spectrum_matter and spectrum_cmb functions select a coarse $k$-grid to solve the perturbations for and interpolates to a finer $k$-grid. This is a common trick to integrate fewer $k$-modes. Precision parameters that affect the output are given with optional keyword arguments. Output spectra and their computation time with standard precision settings are shown in figs. 3 and 1.
It shows that SymBoltz and CLASS agree to around $0.1\%$ or better. The $P(k)$ agree to $0.1\%$ for all $k$ and $0.01\%$ for linear $k$. The $C_{l}$ agree to $0.1\%$ for most $l$, but slightly worse for cosmic variance-dominated $l$ and very large $l$ in $C_{l}^{\text{TE}}$. This is sufficient for most current data (e.g. Bolliet et al., 2024). It is similar to CAMB vs. CLASS with standard precision, although their agreement is pushed to $0.01\%$ with high precision (Lesgourgues, 2011b).
SymBoltz computes $P(k)$ quickly, but the extra steps needed for $C_{l}$ are not yet as optimized as they are in CLASS. A more thorough performance comparison for $P(k)$ is given in section 4.
@@ 3.4 Differentiable Fisher forecasting
Fisher forecasting is a technique to predict how strong parameter constraints that can be placed by data with given errors. It needs derivatives that can be computed with automatic differentiation.
Near a peak $\boldsymbol{p}_{0}$, where derivatives vanish, a log-likelihood $\log L$ of parameters $\boldsymbol{p}$ is approximated by the Taylor series
$\log L(\boldsymbol{p})\approx\log L(\boldsymbol{p}_{0})-\sum_{i,j}F_{ij}(\boldsymbol{p}_{0})(p_{i}-\bar{p}_{i})(p_{j}-\bar{p}_{j}),$ (8)
where $\boldsymbol{F}$ is the Fisher (information) matrix with elements
$F_{ij}(\boldsymbol{p})=-\frac{1}{2}\frac{\partial^{2}\log L(\boldsymbol{p})}{\partial p_{i}\,\partial p_{j}}.$ (9)
Intuitively, $\boldsymbol{F}$ measures how sharp the peak is, or how sensitive $L$ is in different directions in parameter space. Fisher forecasting is powered by the Cramér-Rao bound $\big\lvert\smash{C_{ij}}\big\rvert\geq\big\lvert\smash{F^{-1}_{ij}}\big\rvert$ for the covariance $C_{ij}$ between parameters $p_{i}$ and $p_{j}$. It is an equality for a Gaussian, for which the likelihood (8) is exact. Thus, inverting $\boldsymbol{F}(\boldsymbol{p}_{0})$ gives the tightest parameter constraints one can hope for.
@@
Assuming a $\chi^{2}$-log-likelihood
$\log L(\boldsymbol{p})=-\frac{1}{2}\sum_{l}\bigg(\frac{C_{l}(\boldsymbol{p})-\bar{C}_{l}}{\sigma_{l}}\smash{\bigg)^{2}},$ (11)
the Fisher matrix (9) becomes
$F_{ij}=\sum_{l}\frac{\partial C_{l}}{\partial p_{i}}\frac{1}{\sigma_{l}^{2}}\frac{\partial C_{l}}{\partial p_{j}}.$ (12)
The derivatives $\partial C_{l}/\partial p_{i}$ are usually found with error-prone finite differences by carefully tuning the step. SymBoltz avoids this problem altogether with automatic differentiation: {codebox}
vary = [ M.g.h, M.c.Ω₀, M.b.Ω₀, M.b.YHe, M.ν.Neff, M.h.m_eV, M.X.w₀, M.X.wₐ, M.I.ln_As1e10, M.I.ns,]genprob = parameter_updater(prob, vary)jl, ls = SphericalBesselCache(100:25:1000), 100:1000Cl(p) = spectrum_cmb(:TT, genprob(p), jl, ls)p₀ = map(par -> p[par], vary)dCl_dp = ForwardDiff.jacobian(Cl, p₀)Here vary orders the parameters to differentiate with respect to, and genprob generates a new problem with updated parameters p. This prepares a vector-to-vector function Cl(p) that is differentiated by ForwardDiff.jacobian to compute $\partial C_{l}/\partial p_{i}$ with dual numbers, as shown in fig. 4. Computing and inverting the Fisher matrix (12) forecasts the constraints in fig. 5. They are in excellent agreement with finite difference results from CLASS, but these required significant tuning of precision parameters and step sizes.
Differentiable Fisher forecasts have also been done with an Eisenstein-Hu transfer function fit by Campagne et al. (2023), and through an Einstein-Boltzmann solver by Hahn et al. (2024).
Figure 4: Normalized derivatives $(\partial C_{l}/\partial p_{i})/C_{l}$ of a CMB $T$ power spectrum with respect to cosmological parameters $p_{i}$ from SymBoltz and automatic differentiation (AD; colored lines) versus CLASS and central finite differences (FD; gray dashes). CLASS uses the precision parameters in appendix B and finite differences with $5\%$ relative step sizes.
@@ 3.5 Differentiable MCMC sampling with supernova data
Finally, we show that SymBoltz can output differentiable results for gradient-based MCMC samplers, like the No-U-Turn Sampler (NUTS). It uses gradients to sample parameter space more efficiently than the ?blind? Metropolis-Hastings algorithm.
We predict apparent magnitudes $m(z)=M+\mu(z)$ of Type Ia supernovae at redshift $z$ and fit to 1048 Pantheon observations $(z_{i},m_{i})$ by Scolnic et al. (2018). Their standard absolute magnitude is $M\approx-19.3$, and $\mu(z)=5\lg(d_{L}(z)/10\text{ pc})$ is the distance modulus from the background-derived luminosity distance [^14] This expression is valid for any $\Omega_{k0}$ with complex $\operatorname{sinc}(x)=\sin(x)/x$, but can be split into branches for positive, negative and zero $\Omega_{k0}$.
@@
We define the likelihood $L$ by a multivariate normal distribution of the 1048 supernovae with (constant) covariance matrix $C$ given in detail by Scolnic et al. (2018). To compute $L$, we interface SymBoltz with the probabilistic programming framework Turing.jl (Fjelde et al., 2025), which allows a high-level description of the probabilistic model equivalent to the log-likelihood
$\log L=-\frac{1}{2}\sum_{i,j}C^{-1}_{ij}\big(m(z_{i})-m_{i}\big)\big(m(z_{j})-m_{j}\big).$ (14)
@@
We use the NUTS sampler in Turing. The code for this example can be found in the paper’s notebook. For the flat $w_{0}\text{CDM}$ model ($\Omega_{k0}=0$, $w_{a}=0$, fixed $\Omega_{r0}$) it gives the constraints in fig. 6.
The differentiable computation is not yet fast enough for MCMCs with perturbation-derived spectra. The results here are not groundbreaking, but show that the differentiable pipeline is functional. This is a first step to differentiating the solver quickly, and we aim to accelerate this computation with future work.
Differentiable Boltzmann codes have yet to demonstrate this ability, but more advanced MCMCs with a differentiable background have been done by Campagne et al. (2023), for example.
@@
Figure 6: Parameter inference on 1048 Type Ia supernovae from Pantheon data, using 5000 MCMC samples with the gradient-based NUTS sampler in Turing.jl and differentiable predictions from SymBoltz.
4 Performance comparison to CLASS
Figure 7: Work vs. precision for computing the matter power spectrum $P(k)$ with SymBoltz and CLASS. SymBoltz is approximation-free and run with several implicit ODE solvers. CLASS is run with approximations on/off, its implicit/explicit evolvers ndf15/rk and precision settings in appendix B, but tight coupling is always approximated. For each setup, $P(k)$ is computed by integrating perturbation $k$-modes with ODE solver tolerances $10^{-2}$-$10^{-9}$ that controls the adaptive stepping. For each tolerance, we record the best of 3 runtimes and the $L_{2}$-compressed error $\smash{\big(\sum_{i=1}^{N}\!\big(P(k_{i})/\bar{P}(k_{i})-1\big)^{2}/N\big)^{1/2}}$ relative to high-precision spectra $\bar{P}(k)$ from both codes with approximations off, tolerance $10^{-10}$ and FBDF/ndf15. Both codes solve a $w_{0}w_{a}\text{CDM}$ model with equal parameters, $N=114$ equal wavenumbers $k_{i}$, multipole cutoff $l_{\text{max}}=16$ for all species, default sampling of massive neutrino momenta and parallelize over $k$. Times are from a Linux laptop with Intel i7-12800H CPU.
Figure 7 compares performance versus precision of the matter power spectrum $P(k)$ computed with SymBoltz and CLASS. This computation solves the background, thermodynamics and perturbations for several $k$ and reads off $P(k)$ at the final time. The work is dominated by the perturbations, so it is a relevant test of SymBoltz’ approximation-free treatment of perturbations.
The comparison shows how approximations and the ODE solver affect integration of the Einstein-Boltzmann equations. Without approximations, SymBoltz with Rodas5P is more than $10\times$ faster than CLASS with its implicit ndf15 evolver at comparable precision. When CLASS uses approximations and explicit evolver, SymBoltz is as fast while remaining approximation-free.
To make this comparison as fair as possible, we configured both codes to integrate identical $k$-modes (decided by CLASS) without $k$-interpolation and use the same multipole cutoff $l_{\text{max}}$ for all species. However, the codes sample massive neutrino momenta $q$ differently. This greatly impacts the number of perturbation equations (see section A.7). SymBoltz’ 4 default momenta gives 127 perturbation equations in total. CLASS’ default $\texttt{tol\_ncdm\_newtonian}=10^{-5}$ gives 11 momenta and 245 equations (before approximations). Increasing this tolerance to $10^{-3}$ gives 5 momentum bins, 143 equations and is around $245/143\approx 1.7$ times faster, but introduces errors beyond $1\%$ in $P(k)$. We therefore leave both codes with default $q$-sampling, but note that CLASS could be up to $2\times$ faster at unchanged precision with sparser sampling. This would shift its curves down by $\lg 2\approx 0.3$. Howlett et al. (2012) (in appendix A) also find a precise 3-4 point scheme in CAMB, and Lee et al. (2025) speed up CLASS using integral equations for massive neutrinos in the fork CLASSIER.
Despite lack of approximation schemes, SymBoltz is fast due to high-order implicit solvers like Rodas5P that take accurate and long steps, fast functions for $\boldsymbol{f}$ and the analytically generated $\boldsymbol{J}$, sparse matrix methods with a constant predetermined sparsity structure and efficient sampling of massive neutrino momenta.
SymBoltz is currently slower than CAMB and CLASS for computing $C_{l}$. This is subject to more optimizations than $P(k)$, such as sampling of the source function $S(\tau,k)$, interpolation in $\tau$, $k$, and $l$, evaluation of spherical Bessel functions $j_{l}(x)$ and optimized line-of-sight quadrature. These are very well polished by CAMB and CLASS, and unmatched by SymBoltz at this time. A detailed performance analysis for $C_{l}$ is therefore not meaningful now (but see table 1). With more work, the $C_{l}$ optimizations can be added to SymBoltz. The purpose of the comparison made here is to show the viability of approximation-free perturbations, which was a major milestone in designing SymBoltz.
5 Future work
We hope SymBoltz continues to grow and that others can build on it. Its symbolic capabilities can be extended with utilities for:
  • separation into finer computational stages (see section 2.1.3);
  • gauge transformation of equations (e.g. Newtonian $\leftrightarrow$ synchronous gauge), as already possible in CAMB [^15] https://camb.readthedocs.io/en/latest/symbolic.html;
  • perturbation theory utilities for deriving initial conditions and approximation schemes for new models;
  • transformation of differential equations with respect to conformal time $\tau$ to another (one-to-one) independent variable (e.g. to get observables as direct functions of redshift $z$);
  • validation of dimensions in symbolic equations to catch modeling mistakes by tagging variables with physical units;
  • symbolic tensor manipulation to aid work with modified gravity theories (as a standalone package).
These tasks traditionally involve manual error-prone calculations, so such utilities could ease exploration of extended models.
We particularly aim to improve performance for differentiable runs to the level of standard runs. An attractive milestone is fast reverse-mode gradients of scalar loss functions for gradient-based MCMCs. We anticipate the symbolic structure and Jacobian generation to be useful for this, as writing $\boldsymbol{u}(\tau,\boldsymbol{p})$ in the ODE (3) and differentiating it for $\boldsymbol{v}_{i}=\partial\boldsymbol{u}/\partial p_{i}$ gives $\mathrm{d}\boldsymbol{v}_{i}/\mathrm{d}\tau=\boldsymbol{J}\boldsymbol{v}_{i}+\partial\boldsymbol{f}/\partial p_{i}$, which could be constructed analytically.
It is left for future work to generalize the code from scalar to vector and tensor perturbations, from flat to curved spacetimes, and from adiabatic to other initial conditions. Some additional models like quintessence dark energy and Brans-Dicke gravity are in the works, and we invite others to add extended models.
We hope SymBoltz can foster growth of modular interoperating packages for cosmology in Julia. For example, applications like higher-order perturbation theory (e.g. CLASS-PT by Chudaykin et al. (2020) and PyBird by D’Amico et al. (2021)), nonlinear boosting (e.g. HALOFIT by Takahashi et al. (2012) and HMCODE by Mead et al. (2021)), CMB lensing (e.g. CMBLensing.jl by Millea et al. (2020)) and nonlinear $N$-body simulations all rely on output from Boltzmann solvers. Supporting automatic differentiation across language barriers is hard, and an ecosystem of such packages in one language with support for automatic differentiation, like Julia or Python with JAX, could create a modular and differentiable cosmological modeling toolbox. A recent example of this is DISCO-DJ’s coupling of differentiable Boltzmann and $N$-body codes for end-to-end modeling from linear perturbations to non-linear structure formation (Hahn et al., 2024; List et al., 2025).
6 Conclusion